# Repositioning, Ride-matching, and Abandonment in On-demand Ride-hailing Platforms: A Mean Field Game Approach

### v1

@@ 4. Dynamically Optimize Matching Radii
In this section, we demonstrate how to address the “Wild Goose Chase” problem and eliminate multiple equilibria by implementing an optimal dynamic adjustment of matching radii.
The multiple equilibria and WGC are not resolved by introducing fixed matching radii. Instead of treating matching radii as fixed parameters of the dispatch algorithm, we propose to dynamically optimize matching radii. Specifically, we treat the waiting time $w_{i}^{d}$ and pick-up time $\tau_{i}$ in (21) as functions of both the driver supply rate $x_{ii}$ and the matching radii $R_{i,c}$ and $R_{i,d}$. Our goal is to find the optimal radii $R_{i,c}$ and $R_{i,d}$ that minimize the sum of waiting and pick-up times:
(25) $(R_{i,c}(x_{ii}),R_{i,c}(x_{ii}))=\arg\min_{(R_{i,c},R_{i,d})\in\mathbb{R}^{2}_{+}}w_{i}^{d}(x_{ii},R_{i,c},R_{i,d})+\tau_{i}(x_{ii},R_{i,c},R_{i,d}),\quad\forall i\in S,$
where $w_{i}^{d}$ and $\tau_{i}$ are given by approximation formulas in (21) since the original formulas (18) and (19) we derive can be approximated by (21) when $R_{c},R_{d}$ are small. We will show that the above optimization problem in (25) can be solved and the dynamic matching radii function $(R_{i,c}(x_{ii}),R_{i,d}(x_{ii}))$ is well-defined as the unique minimizer of $w_{i}^{d}+\tau_{i}$ for each $x_{ii}$ in the following proposition.
@@
Proof.
The optimization problem (26) is equivalent to the following problem,
s.t. $\quad x_{ii}e^{-\alpha_{i}^{2}}+b_{i}e^{-\beta_{i}^{2}}=b_{i}$
over $\quad u_{i},\alpha_{i},\beta_{i}\in\mathbb{R}_{+}$
which is obtained from (26) by changes of variables as follows,
[…]
In the following theorem, we demonstrate that the sojourn time $T_{ii}(x_{ii})$ as defined in Assumption 1 is indeed a strictly increasing function of $x_{ii}$ for all $i\in\mathcal{S}$ if the proposed NN dispatch algorithm with dynamic radii is adopted.
@@
Given the optimal adjustment of matching radii $(R_{i,c}(x_{ii}),R_{i,d}(x_{ii}))$ defined by (25), i.e., $R_{i,c}(x_{ii})=R_{i,d}(x_{ii})=R_{i}^{*}(x_{ii})$, where $R_{i}^{*}(x_{ii})$ is the unique solution to (27), the following function
@@
is a strictly increasing function of $x_{ii}$ when $x_{ii}\in(0,b_{i})$ for all $i\in\mathcal{S}$. Consequently, there exists a unqiue stationary equilibrium for game $\mathcal{G}$.
Proof.
where
and $R_{i}^{*}(x_{ii})$ is the unique solution to (27). One can check that this derivative is strictly positive when $x_{ii}\in(0,b_{i})$ for all $i\in\mathcal{S}$. Then, the uniqueness of stationary equilibrium follows from Corollary 2.5. ∎
@@ 5. Simulation
We have theoretically proved and numerically shown that the proposed optimal adjustment of matching radius resolves the WGC problem and ensures a unique equilibrium. To further verify the effectiveness of the proposed method, we simulate the stochastic online matching process in a single region. We apply the nearest-neighbor dispatch with the optimal adjustment of matching radius where the optimal matching radius is computed for each value of driver supply rate using the equation (27). The experiments are performed on using Python 3.11 on a 16-core CPU.
@@ 6. Conclusion
In this study, we have made significant strides in understanding and enhancing the operational dynamics of ride-hailing platforms through the lens of mean-field game theory. Our primary contribution lies in the introduction of a novel mean field game model, which not only systematically analyzes the intricate dynamics of these platforms but also establishes a profound connection between the equilibrium points of the game and the KKT points of an associated optimization problem. This theoretical framework offers a fresh perspective on platform operations, enabling a deeper understanding of the underlying mechanisms that govern driver and customer interactions. Furthermore, our analysis of ride-matching algorithms in the presence of customer abandonment behavior has unveiled critical insights. We have identified that these algorithms can inadvertently lead to non-monotone sojourn times, resulting in the emergence of multiple equilibria and the problematic “Wild Goose Chase” one. This WGC equilibrium, characterized by drivers engaging in inefficient pursuits of distant requests, poses a significant challenge to the overall efficiency of ride-hailing systems. To address these inefficiencies, we have proposed an innovative two-matching-radius nearest-neighbor dispatch algorithm. Our rigorous theoretical proofs and empirical validations demonstrate that the dynamic and optimal adjustment of matching radii can effectively ensure that sojourn times increase monotonically with the driver supply rate. This breakthrough not only provides a robust solution to the WGC problem but also enhances the overall efficiency and reliability of ride-hailing platforms.
@@
We conclude this paper by envisioning a roadmap for future research that extends the boundaries of our current work. One promising avenue lies in the exploration of pricing mechanisms within ride-hailing platforms. Inspired by (Castillo et al., 2023)’s findings, which indicate that surge pricing outperforms matching adjustments in mitigating the WGC equilibrium, we aim to conduct a comprehensive and theoretical comparison using our mean field game framework. The findings in current work suggest that an optimal matching strategy can effectively eliminate WGC without reliance on any pricing designs. We expect that incorporating pricing design into our model will unveil new dimensions and insights, further enriching our understanding of ride-hailing dynamics. Another extension of our model is its application to delivery problems, where couriers, analogous to drivers in our setting, have the capability to deliver multiple parcels simultaneously. The analysis of spatial queueing networks in this context is more complex and challenging. We are particularly interested in exploring whether our approach of expressing sojourn times as functions of throughput can be adapted and applied to this more intricate setting. Such an extension could provide valuable insights into the optimization of delivery systems, contributing to enhanced efficiency and service quality.

### v2

@@ 4. Dynamically Optimize Matching Radii
In this section, we demonstrate how to eliminate multiple equilibria by implementing an optimal adjustment of matching radii.
The multiple equilibria are not resolved by introducing fixed matching radii. Instead of treating matching radii as fixed parameters of the dispatch algorithm, we propose to dynamically optimize them. Specifically, we treat the waiting time $w_{i}^{d}$ and pick-up time $\tau_{i}$ in (19) as functions of both the driver supply rate $x_{ii}$ and the matching radii $R_{i,c}$ and $R_{i,d}$, and use the optimal radii $R_{i,c}$ and $R_{i,d}$ that minimize the sum of waiting and pick-up times:
(24) $(R_{i,c}(x_{ii}),R_{i,c}(x_{ii}))=\arg\min_{(R_{i,c},R_{i,d})\in\mathbb{R}^{2}_{+}}w_{i}^{d}(x_{ii},R_{i,c},R_{i,d})+\tau_{i}(x_{ii},R_{i,c},R_{i,d}),\quad\forall i\in S.$
Here, $w_{i}^{d}$ and $\tau_{i}$ are given by approximation formulas in (19) since the original formulas (16) and (17) can be approximated by (19) when $R_{c},R_{d}$ are small. We show that the above optimization problem in (24) can be solved analytically and the dynamic matching radii function $(R_{i,c}(x_{ii}),R_{i,d}(x_{ii}))$ is well-defined as the unique minimizer of $w_{i}^{d}+\tau_{i}$ for each $x_{ii}$, in the following proposition.
@@
The proof can be found in Appendix C. Proposition 4.1 show that for any driver supply rate $x_{ii}>0$ in any region $i\in S$, the optimal choice of matching radii is unique and can be found by solving (26). In the following theorem, we demonstrate that the sojourn time $T_{ii}(x_{ii})$ as defined in Assumption 1 is indeed a strictly increasing function of $x_{ii}$ for all $i\in\mathcal{S}$ if the proposed dispatch algorithm with optimal adjustment of matching radii in (24) is adopted.
@@
Given the optimal adjustment of matching radii $(R_{i,c}(x_{ii}),R_{i,d}(x_{ii}))$ defined by (24), i.e., $R_{i,c}(x_{ii})=R_{i,d}(x_{ii})=R_{i}^{*}(x_{ii})$, where $R_{i}^{*}(x_{ii})$ is the unique solution to (26), the function
@@
is a strictly increasing in $x_{ii}$ when $x_{ii}\in(0,b_{i})$ for all $i\in\mathcal{S}$. Consequently, there exists a unique stationary equilibrium for game $\mathcal{G}$.
The proof can be found in Appendix D. Theorem 4.2 establishes that the optimal adjustment of matching radii, as defined in (24), ensures a unique equilibrium, thereby eliminating the inefficient equilibrium with long pick-up times. Importantly, this theorem also demonstrates that our optimal adjustment guarantees the optimization problem (10) is strictly convex–a property whose significance has not been sufficiently emphasized so far. When the optimization problem (10) is nonconvex, multiple equilibria may arise, and computing these equilibria (or KKT points) becomes computationally challenging. In contrast, the strict convexity of (10) enables the use of efficient gradient-based numerical methods to compute the unique equilibrium, enhancing the stability and predictability of the ride-hailing system. Moreover, by minimizing the driver’s total waiting time (the sum of waiting and pick-up times) in each region, our proposed algorithm also enhances the efficiency of the ride-hailing system. The improvement is further confirmed through the numerical experiments and simulations conducted for a single region.
@@ 5. Simulation
We have theoretically proved and numerically shown that the proposed optimal adjustment of matching radius ensures a unique equilibrium and improve efficiency. To further verify the effectiveness of the proposed method, we simulate the stochastic online matching process in a single region. We apply the nearest-neighbor dispatch with the optimal adjustment of matching radius where the optimal matching radius is computed for each value of driver supply rate using the equation (26). The experiments are performed using Python 3.11 on a 16-core CPU.
@@ 6. Conclusion
In this study, we have made significant strides in understanding and enhancing the operational dynamics of ride-hailing platforms through the lens of mean-field game theory. Our primary contribution lies in the introduction of a novel mean field game model, which not only systematically analyzes the intricate dynamics of these platforms but also establishes a profound connection between the equilibrium points of the game and the KKT points of an associated optimization problem. This theoretical framework offers a fresh perspective on platform operations, enabling a deeper understanding of the underlying mechanisms that govern driver and customer interactions. Furthermore, our analysis of ride-matching algorithms in the presence of customer abandonment behavior has unveiled critical insights. We have identified that these algorithms can inadvertently lead to non-monotone sojourn times, resulting in the emergence of multiple equilibria and the problematic inefficient ones. The inefficient equilibria, characterized by drivers engaging in inefficient pursuits of distant requests, pose a significant challenge to the overall efficiency of ride-hailing systems. To address these inefficiencies, we have proposed an innovative two-matching-radius nearest-neighbor dispatch algorithm. Our rigorous theoretical proofs and empirical validations demonstrate that the dynamic and optimal adjustment of matching radii can effectively ensure that sojourn times increase monotonically with the driver supply rate. This breakthrough not only provides a robust solution to the multiple and inefficient equilibria problem but also enhances the overall efficiency and reliability of ride-hailing platforms.
@@
We conclude this paper by envisioning a roadmap for future research that extends the boundaries of our current work. One promising avenue lies in the exploration of pricing mechanisms within ride-hailing platforms. Inspired by (Castillo et al., 2023)’s findings, which indicate that surge pricing outperforms matching adjustments in mitigating the inefficient equilibria, we aim to conduct a comprehensive and theoretical comparison using our mean field game framework. The findings in current work suggest that an optimal matching strategy can effectively eliminate inefficient equilibria without reliance on any pricing designs. We expect that incorporating pricing design into our model will unveil new dimensions and insights, further enriching our understanding of ride-hailing dynamics. Another extension of our model is its application to delivery problems, where couriers, analogous to drivers in our setting, have the capability to deliver multiple parcels simultaneously. The analysis of spatial queueing networks in this context is more complex and challenging. We are particularly interested in exploring whether our approach of expressing sojourn times as functions of throughput can be adapted and applied to this more intricate setting. Such an extension could provide valuable insights into the optimization of delivery systems, contributing to enhanced efficiency and service quality.
Appendix A Derivation of Formulas for Waiting Times and Pick-up Times
In this appendix we establish formulas for the average waiting time and pick-up time as claimed in Assumption 1 based on the customer abandonment behavior defined in Section 3.1 and the platform’s dispatching algorithm defined in Section 3.2.
Let each region $i$’s area be denoted by $a_{i}$. We introduce two matching radii in region $i$, $R_{i,c},R_{i,d}\in(0,\sqrt{a_{i}/\pi})$, where $\sqrt{a_{i}/\pi}$ is the approximate radius of region $i$. We denote the average vehicle speed by $v_{i}$. Let $\mu_{i}^{c}$ and $\mu_{i}^{d}$ denote the average densities of waiting customers and waiting drivers respectively across the region $i$. Recall that $x_{ii}$ denotes the rate of drivers that become available and choose to serve region $i$ per unit area, and $b_{i}$, $\theta_{i}$ the arrival rate, abandonment rate of customers per unit area, respectively.
$(1-\frac{\pi R_{i,c}^{2}}{a_{i}})^{N_{i}^{c}}.$
Similarly, given the matching radius $R_{i,d}$, if there are $N_{i}^{d}$ waiting drivers in region $i$, a new arriving customer is not matched with any of these waiting customers with probability
$(1-\frac{\pi R_{i,d}^{2}}{a_{i}})^{N_{i}^{d}}.$
(27) $x_{ii}(1-\frac{\pi R_{i,c}^{2}}{a_{i}})^{\mu_{i}^{c}a_{i}}+b_{i}(1-\frac{\pi R_{i,d}^{2}}{a_{i}})^{\mu_{i}^{d}a_{i}}=b_{i}\,.$
(28) $\mu_{i}^{c}(x_{ii})=\frac{1}{\theta}(b_{i}-x_{ii}).$
As a result, by (15), the unknown variable $\mu_{i}^{d}$ can also be expressed as a function of $x_{ii}$,
Next, we derive the formula for expected pick-up time, which amounts to finding the expected pick-up distance. We compute this in two distinct cases:
    $1-(1-\frac{\pi R_{i,c}^{2}}{a_{i}})^{\mu_{i}^{c}a_{i}}.$
    $1-(1-\frac{\pi r^{2}}{a_{i}})^{\mu_{i}^{c}a_{i}}\,.$
    $(1-\frac{\pi R_{i,c}^{2}}{a_{i}})^{\mu_{i}^{c}a_{i}}.$
    $1-(1-\frac{\pi R_{i,d}^{2}}{a_{i}})^{\mu_{i}^{d}a_{i}}.$
    $1-(1-\frac{\pi r^{2}}{a_{i}})^{\mu_{i}^{d}a_{i}}\,.$
In summary, the expected pick-up time is a function of $x_{ii}$ as follows,
where $\mu_{i}^{c}$ and $\mu_{i}^{d}$ are functions of $x_{ii}$ as in (LABEL:eq:u_c) and (29) respectively.
$x_{ii}e^{-\mu_{i}^{c}\pi R_{i,c}^{2}}+b_{i}e^{-\mu_{i}^{d}\pi R_{i,d}^{2}}=b_{i}.$
(33) $\begin{cases}w_{i}^{d}(x_{ii})&=\frac{-\log(1-\frac{x_{ii}}{b_{i}}e^{-\frac{b_{i}-x_{ii}}{\theta_{i}}\pi R_{i,c}^{2}})}{x_{ii}\pi R_{i,d}^{2}}\,,\\ \tau_{i}(x_{ii})&=\frac{1}{v_{i}}\int_{0}^{R_{i,c}}e^{-\mu_{i}^{c}(x_{ii})\pi r^{2}}-e^{-\mu_{i}^{c}(x_{ii})\pi R_{i,c}^{2}}\mathrm{d}r+\frac{b_{i}}{v_{i}x_{ii}}\int_{0}^{R_{i,d}}e^{-\mu_{i}^{d}(x_{ii})\pi r^{2}}-e^{-\mu_{i}^{d}(x_{ii})\pi R_{i,d}^{2}}\mathrm{d}r\,.\end{cases}$
Note that the left-hand side of (18) is approximately $R_{i,\cdot}^{2}$ when $R_{i,\cdot}<<\sqrt{a_{i}}$ as $\log(1-y)\approx-y$ when $y$ approaches 0 and goes to $\infty$ as $\pi R_{i,\cdot}^{2}\rightarrow a_{i}$. Hence, when calculating the pick-up time we do not multiply the change of variable factor when computing the integral. Numerical results show that this is still a good estimation of the original pick-up time even when the original radius $R_{i}$ is large. There is an alternative point of view which yields the same approximate formula in (19): assume each region $i$ is an unbounded plane with matching radii $R_{i,c},R_{i,d}\in(0,\infty)$. In this unbounded plane, the locations of waiting customers and drivers follow spatial Poisson processes, where $\mu_{i}^{c}$ and $\mu_{i}^{d}$ are the waiting customer and waiting driver intensity parameters, respectively. In fact, the change of variable in equation (18) defines a map from the original finite region to an infinite region. Although this transformation distorts the Euclidean metric in the original region, the distortion is negligible when the matching radius is sufficiently small. For large matching radii, while the distortion cannot be entirely ignored, the introduced approximation error remains acceptable according to numerical results.
Appendix B Approximate Pick-up Time by Square Root Law
When $R_{c}=R_{d}=\sqrt{a/\pi}$ (or $\infty$), our formulae for $\tau_{i}$ in (17) and (19) can still be used to estimate the pick-up time. In fact, both formulas give the same asymptotic result. We assume $\mu^{c}=0,\mu^{d}>0$. In this case, $x=b$ as $\mu^{c}=0$. Applying (17), the average pick-up time is estimated by
Appendix C Proof of Proposition 4.1
Proof.
The optimization problem (25) is equivalent to the following problem,
s.t. $\quad x_{ii}e^{-\alpha_{i}^{2}}+b_{i}e^{-\beta_{i}^{2}}=b_{i}$
over $\quad u_{i},\alpha_{i},\beta_{i}\in\mathbb{R}_{+}$
which is obtained from (25) by changes of variables as follows,
[…]
Appendix D Proof of Theorem 4.2
Proof.
where
and $R_{i}^{*}(x_{ii})$ is the unique solution to (26). One can check that this derivative is strictly positive when $x_{ii}\in(0,b_{i})$ for all $i\in\mathcal{S}$. Since $f_{i}(x_{ii})$ is a strictly increasing function of $x_{ii}$ when $x_{ii}\in(0,b_{i})$ for all $i\in\mathcal{S}$, the uniqueness of stationary equilibrium follows from Corollary 2.5. ∎